The treasurer of a Canadian company has C$1,000,000 to invest for 90 days.  A 90-day U.S. commercial paper offers a yield of 5.00 percent.  The present exchange rate of C$ is $0.8980.  If the Canadian company decided to invest in U.S. commercial paper, what is the expected (annual) yield if the expected exchange rate of CS at the end of 90 days is $0.8890?  (You will find an illustrative example for this type questions in the course content folder)
Answer #1

Spot Exchange Rate (C$/$) = 0.8980

Converting C$1000000 into $

1000000*0.8980 = $898000

Invest it at 5% yield for 90 days

Amount after 90 days = 898000(1+(0.05/365))90 = $909139

Converting $909139 into C$

$909139/0.8890 = C$1022653.54

Yield =(1022653.54-1000000)/1000000 = 2.265%

Annual Yield = 2.265*365/90 = 9.187%

*Assuming 1 year = 365 days